OpenStudy (anonymous):

What value of k solves the equation? k^-4= 1/256 A. –64 B. 8 C. 4 D. 64

OpenStudy (anonymous):

@imqwerty

imqwerty (imqwerty):

\[256=4^4\]and\[\frac{ 1 }{ 256 }=\frac{ 1 }{ 4^4 }=4^\left( -4 \right)\]but we are given that\[k^\left( -4 \right)=\frac{ 1 }{ 256 }\]and \[\frac{ 1 }{ 256 }=4^\left( -4 \right)\]so\[k^\left( -4 \right)=4^\left( -4 \right)\]comparing u get k=4